# Energy-Efficient Multi-UAV-Enabled MEC Systems over Space-Air-Ground Integrated Networks

### v2

@@ III-C Computing Model
where $\kappa_{\rm{L}}$ denotes the energy efficiency factor for the local computing of all MUs [36].
UAV Processing
When the MU $m$ offloads a partial task $\omega_{m}^{\rm{U}}[n]D_{m}[n]$ to the associated UAV, the latency comprises three parts: the time for task data uplink transmission from the MU to the associated UAV, the time for the task computation at the associated UAV, and the time for task computation result downlink transmission from the UAV to the MU. Since the data volume of the results is much smaller than that of the offloaded task, the download latency can be neglected [11].
@@
where $\kappa_{\rm{U}}$ denotes the energy efficiency factor for task processing of the UAVs [36].
LEO Satellite Processing
@@
where $\kappa_{\rm{S}}$ denotes the energy efficiency factor for task computing of the LEO satellite [36].
@@ III-D Problem Formulation
In this work, we aim to maximize the total energy efficiency of the system by jointly optimizing the UAV trajectory $\boldsymbol{Q}=\{\boldsymbol{q}_{k}[n],\forall k,n\}$, the transmit power of all MUs $\boldsymbol{P}=\{p_{m,k}[n],p_{m,{\rm{LEO}}}[n],\forall m,k,n\}$, the task offloading decision of all MUs $\boldsymbol{\Omega}=\{\omega_{m}^{\rm{L}}[n],\omega_{m}^{\rm{U}}[n],\omega_{m}^{\rm{S}}[n],\forall m,n\}$, the association among the MUs and the UAVs $\boldsymbol{\alpha}=\{\alpha_{m,k}[n],\forall m,k,n\}$, and the computing frequency allocation of the MUs, the UAVs and the LEO satellite $\boldsymbol{F}=\{f^{\rm{L}}_{m}[n],f^{\rm{U}}_{m,k}[n],f^{\rm{S}}_{m}[n],\forall m,k,n\}$. The problem can be formulated as 111 Noting that the proposed SAGIN-supported MEC system in this paper can be directly extended to the case with varying numbers of UAVs and MUs, when the associations of the MUs changes. Also the system can be extended to the case with dynamic environments when the channel has different modeling.
@@
${\rm{s.t.}}$ $\boldsymbol{q}_{k}[1]=\boldsymbol{q}_{k,\mathsf{I}},\boldsymbol{q}_{k}[N]=\boldsymbol{q}_{k,\mathsf{F}},\forall k,$ (17b)
@@
where $F^{\rm{L}}_{\rm{max}}$, $F^{\rm{U}}_{\rm{max}}$, and $F^{\rm{S}}_{\rm{max}}$ are the maximum computing frequencies of the MUs, the UAVs, and the LEO satellite, respectively. $P_{\rm{max}}^{\rm{U}}$ and $P_{\rm{max}}^{\rm{S}}$ denote the maximum transmit power of the MUs for communicating with the associated UAV, and the LEO satellite, respectively. $E_{\rm{max}}^{{\rm{UAV}}}$ denotes the energy budget of the UAVs. The constraints (17b)-(17d) represent the mobility constraints of the UAVs. The constraints (17e)-(17f) represent the association constraints among the MUs and the UAVs. The constraints (17g)-(17h) represent the task offloading decision constraints. The constraints (17i)-(17k) represent the constraints on the computing frequencies for local task processing, the computing frequencies for the UAV task processing, and the computing frequencies for the LEO satellite task processing. The constraints (17l)-(17m) represent the transmit power allocation from the MUs to the UAVs, and the transmit power allocation from the MUs to the LEO satellite. The constraint (17n) represents the latency constraint. Lastly, the constraint (17o) represents the constraint on the energy consumption of the UAVs 222 In larger-scale scenarios, the energy consumption of UAVs emerges as a critical bottleneck, stemming from extended flight distances and escalating task computational overhead. Battery limitations would require careful consideration of energy efficient strategies. We will consider novel battery models, including energy harvesting and periodic recharging, to improve the energy-efficiency of the system in our future works. .
@@ IV-A Quadratic Transformation-based Problem Reformulation
Please refer to [48]. ∎
@@ IV-B Subproblem 1: MU-UAV Association Optimization
Problem SP1 is a standard integer linear programming problem, which can be solved using conventional optimization methods, such as the branch-and-bound algorithm [49].
@@ IV-C Subproblem 2: Task Offloading Decision Optimization
Problem SP2 is a standard linear programming problem, which can be solved using the interior-point method [50].
@@ IV-E Subproblem 4: UAV Trajectory Optimization
$\begin{split}&\psi_{k}^{2}[n]\tau^{2}\leq-\|\boldsymbol{q}_{k}^{(l)}-\boldsymbol{q}_{k}^{(l)}[n-1]\|^{2}\\ &+2(\boldsymbol{q}_{k}^{(l)}[n]-\boldsymbol{q}_{k}^{(l)}[n-1])^{\mathsf{T}}(\boldsymbol{q}_{k}[n]-\boldsymbol{q}_{k}[n-1]).\end{split}$ (45)
@@
IV-F Overall Algorithm and Complexity Analysis
@@ IV-F2 Complexity Anlysis
In Algorithm 1, we solve the subproblem SP1 using the binary cut-and-branch method [49], and solve other subproblems using the interior point method [50]. Thus, the computational complexity of the four subproblems are $\mathcal{O}\left(MKN\log(MKN)\right)$, $\mathcal{O}\left((MN)^{3.5}\log(1/\epsilon)\right)$, $\mathcal{O}\left((MKN)^{3.5}\log(1/\epsilon)\right)$, $\mathcal{O}\left((MKN)^{3.5}\log(1/\epsilon)\right)$, respectively, where $\epsilon$ represents the tolerance. Consequently, the complexity of Algorithm 1 is $\mathcal{O}(\mathcal{I}(((MN)^{3.5}$ $+(MKN)^{3.5}+(MKN)^{3.5})\log(1/\epsilon)+MKN\log(MKN)))$, where $\mathcal{I}$ denotes the iteration number.
@@ V Numerical Results
In the simulations, we consider $K=2$ UAVs fly from specified initial positions to specified final positions to provide computing services for the MUs. We consider $M=8$ MUs that are randomly distributed in a square area of $1000\times 1000\ \rm{m}^{2}$. All UAVs fly at a height of 100 $\rm{m}$. Unless otherwise stated, the remaining simulation parameters are summarized in Table I [52, 36].
TABLE I: Simulation Parameters.
@@
Number of time slot, $N$ 40
@@
Channel power gain between the UAVs and the MUs, $\beta_{0}$60 dBm
Noise power at the UAV, $\sigma_{\rm{UAV}}$170 dBm
@@
Noise power at the LEO satellite, $\sigma_{\rm{LEO}}$170 dBm
@@
To evaluate the performance and efficiency of the proposed scheme, we compare it with five baseline schemes, which are given in the following.
@@
    LEO satellite (LS) Offloading Scheme: Only the LEO satellite provides computing services for the MUs.
  • 5)
    Hybrid LEO satellite and BS (HLAB) Offloading Scheme: The LEO Satellite and BS can provide computing services for the MUs in hybrid mode.
@@
Fig. 5 and Fig. 6 show that, as the number of MUs increases, our proposed scheme exhibits a more significant increase in total energy efficiency compared to other five schemes, while the increase in energy consumption is relatively smaller. The reason can be explained as follows. As the number of MUs grows, the demand for the UAV services increases, and the various schemes exhibit different levels of adaptability to these changes. Specifically, the HLAB’s offloading capability is confined to the BS and the LEO satellite while neglecting the UAV mobility advantages, resulting in lower efficiency of the air-to-ground communication link. The single UAV scheme deploys only a single UAV, which makes it challenging to ensure high-quality communication services for every MU when a large number of MUs require for the services. The fixed trajectory scheme lacks flexibility, especially when there are multiple UAVs. This means that in scenarios with the uneven MU distribution or dynamic demand changes, the scheme struggles to optimize service quality by adjusting the UAV positions. The fixed allocation scheme employs a fixed task allocation decision, which means it cannot adjust the allocation of tasks based on real-time network conditions. In contrast, our scheme dynamically reallocates tasks according to the current network state, ensuring higher service quality and better user experience. The LS offloading scheme relies solely on local computing resources and the LEO satellite assisted computing without considering the role of the UAVs. Although this approach can provide a certain level of service, the absence of the UAV support makes it difficult to achieve ideal computational quality and efficiency in scenarios with a large-scale MU distribution. Therefore, under the same number of MUs, our proposed scheme demonstrates the best performance. Specifically, from the numerical results, when the number of the MUs is $16$, our proposed scheme outperforms the HLAB offloading scheme, the single UAV scheme, fixed trajectory scheme, fixed allocation scheme, and LS offloading scheme by $13.74\%$, $17.53\%$, $22.55\%$, $48.27\%$, and $133.90\%$, respectively, in terms of the overall computational energy efficiency.
@@
Fig. 7 and Fig. 8 show that, as the amount of task data that the MUs need to process per time slot increases, the total energy efficiency of all schemes decreases, and the required energy consumption also increases. Notably, in both key metrics of total energy efficiency and energy consumption, our proposed algorithm outperforms other five algorithms. The reason is that as the amount of task data increases, the MUs tend to offload more tasks to the UAVs and the LEO satellite for remote computing, leading to an increase in communication energy consumption, which in turn reduces the total energy efficiency of all schemes. Moreover, the LS offloading scheme experiences the most significant increase in total energy consumption, and the rate of increase is notably faster than that of other five schemes. The reason is that, in the LS offloading scheme, the MUs do not offload tasks to the UAVs, which concentrate data processing burdens exclusively on the MUs and the LEO satellite.
@@
Fig. 9 presents the total computational energy efficiency versus the communication bandwidth between the UAVs and the MUs. It can be observed that the total energy efficiency significantly increases with the growth of the communication bandwidth between the MUs and the UAVs for the proposed scheme, the HLAB offloading scheme (noting that in HLAB offloading scheme, the MU-UAV links are replaced with the MU-BS links), single UAV scheme, and fixed trajectory scheme, whereas the increase is relatively slower for fixed allocation scheme, and remains unchanged for LS offloading scheme. The reason is that the increase in communication bandwidth enhances the task transmission rate between the MUs and the UAVs, allowing the UAVs to assist in processing a larger volume of tasks within each time slot $\tau$. Specifically, for the proposed scheme, the HLAB offloading scheme, single UAV scheme, and fixed trajectory scheme, the increase in communication bandwidth means that the MUs can offload tasks to the UAVs more quickly, improving the efficiency of task processing and reducing the burden of local computation, thus enhancing overall energy efficiency. However, in the fixed allocation scheme, since the proportions of task offloading to the UAVs is fixed, the growth in communication bandwidth can only reduce the energy consumption associated with the UAV computing, but does not enable the MUs to offload more tasks to the UAV, thereby limiting its potential for improvement in energy efficiency. For the LS offloading scheme, since there is no UAV involved, the increase in communication bandwidth has no effect, and the total energy efficiency remains a constant.
Fig. 10 presents the total energy consumption versus the communication bandwidth between the UAVs and the MUs. The proposed scheme, the HLAB offloading scheme, single UAV scheme, and fixed trajectory scheme demonstrate a significant reduction in total energy consumption as the communication bandwidth increases, whereas the reduction is relatively slower for the fixed allocation scheme, and remains unchanged for the LS offloading scheme. The reason is that the increase in communication bandwidth significantly enhances the communication rate between the MUs and the UAVs, thereby increasing the proportion of tasks offloaded from the MUs to the UAVs, which is shown in Fig. 11. Specifically, when the communication bandwidth increases from $1$ MHz to $2$ MHz, the average proportion of tasks offloaded to the UAVs rises from $49\%$ to $65\%$, while the average proportion of tasks processed locally decreases from $31\%$ to $24\%$ in the proposed scheme. Since the MUs are equipped with low energy-efficient processors, offloading more tasks to the UAVs for processing can effectively lower the total energy consumption of the system.
@@
Fig. 13 illustrates the total energy efficiency under different schemes versus the maximum tolerable latency per time slot. Here, the maximum tolerable latency corresponds to the time slot duration $\tau$. It can be observed that as the maximum tolerable latency increases, all schemes exhibit monotonically increasing energy efficiency due to the extended temporal flexibility for data offloading and processing. Notably, our proposed scheme consistently maintains the highest performance across all the maximum tolerable latency regimes, stemming from the synergistic data processing coordination among the MUs, the UAVs, and the LEO satellite.
Figure 13: The total energy efficiency under different schemes versus the maximum tolerable latency per time slot.

### v3

@@ III-C1 Local Processing
where $\kappa_{\rm{L}}$ denotes the energy efficiency factor for the local computing of all MUs.
III-C2 UAV Processing
When the MU $m$ offloads a partial task $\omega_{m}^{\rm{U}}[n]D_{m}[n]$ to the associated UAV, the latency comprises three parts: the time for task data uplink transmission from the MU to the associated UAV, the time for the task computation at the associated UAV, and the time for task computation result downlink transmission from the UAV to the MU. Since the data volume of the results is much smaller than that of the offloaded task, the download latency can be neglected [Liu2023Energy-Efficient].
@@ III-C2 UAV Processing
where $\kappa_{\rm{U}}$ denotes the energy efficiency factor for task processing of the UAVs.
III-C3 LEO Satellite Processing
@@
where $\kappa_{\rm{S}}$ denotes the energy efficiency factor for task computing of the LEO satellite.
@@ III-D PROBLEM FORMULATION
In this work, we aim to maximize the total energy efficiency of the system by jointly optimizing the UAV trajectory $\boldsymbol{Q}=\{\boldsymbol{q}_{k}[n],\forall k,n\}$, the transmit power of all MUs $\boldsymbol{P}=\{p_{m,k}[n],p_{m,{\rm{LEO}}}[n],\forall m,k,n\}$, the task offloading decision of all MUs $\boldsymbol{\Omega}=\{\omega_{m}^{\rm{L}}[n],\omega_{m}^{\rm{U}}[n],\omega_{m}^{\rm{S}}[n],\forall m,n\}$, the association among the MUs and the UAVs $\boldsymbol{\alpha}=\{\alpha_{m,k}[n],\forall m,k,n\}$, and the computing frequency allocation of the MUs, the UAVs and the LEO satellite $\boldsymbol{F}=\{f^{\rm{L}}_{m}[n],f^{\rm{U}}_{m,k}[n],f^{\rm{S}}_{m}[n],\forall m,k,n\}$. The problem can be formulated as
@@
${\rm{s.t.}}\$$\boldsymbol{q}_{k}[1]=\boldsymbol{q}_{k,\mathsf{I}},\boldsymbol{q}_{k}[N]=\boldsymbol{q}_{k,\mathsf{F}},\forall k,$ (17b)
@@
where $F^{\rm{L}}_{\rm{max}}$, $F^{\rm{U}}_{\rm{max}}$, and $F^{\rm{S}}_{\rm{max}}$ are the maximum computing frequencies of the MUs, the UAVs, and the LEO satellite, respectively. $P_{\rm{max}}^{\rm{U}}$ and $P_{\rm{max}}^{\rm{S}}$ denote the maximum transmit power of the MUs for communicating with the associated UAV, and the LEO satellite, respectively. $E_{\rm{max}}^{{\rm{UAV}}}$ denotes the energy budget of the UAVs. The constraints (17b)-(17d) represent the mobility constraints of the UAVs. The constraints (17e)-(17f) represent the association constraints among the MUs and the UAVs. The constraints (17g)-(17h) represent the task offloading decision constraints. The constraints (17i)-(17k) represent the constraints on the computing frequencies for local task processing, the computing frequencies for the UAV task processing, and the computing frequencies for the LEO satellite task processing. The constraints (17l)-(17m) represent the transmit power allocation from the MUs to the UAVs, and the transmit power allocation from the MUs to the LEO satellite. The constraint (17n) represents the latency constraint. Lastly, the constraint (17o) represents the constraint on the energy consumption of the UAVs.
@@ IV-A QUADRATIC TRANSFORMATION-BASED PROBLEM REFORMULATION
Please refer to [Shen2018Fractional1]. ∎
@@ IV-B SUBPROBLEM 1: MU-UAV ASSOCIATION OPTIMIZATION
Problem SP1 is a standard integer linear programming problem, which can be solved using conventional optimization methods, such as the branch-and-bound algorithm [2013MunozJoint].
@@ IV-C SUBPROBLEM 2: TASK OFFLOADING DECISION OPTIMIZATION
Problem SP2 is a standard linear programming problem, which can be solved using the interior-point method [2010LuoSemidefinite].
@@ IV-E SUBPROBLEM 4: UAV TRAJECTORY OPTIMIZATION
$\begin{split}&\psi_{k}^{2}[n]\tau^{2}\leq\|\boldsymbol{q}_{k}^{(l)}-\boldsymbol{q}_{k}^{(l)}[n-1]\|^{2}\\ &+2(\boldsymbol{q}_{k}^{(l)}[n]-\boldsymbol{q}_{k}^{(l)}[n-1])^{\mathsf{T}}(\boldsymbol{q}_{k}[n]-\boldsymbol{q}_{k}[n-1]).\end{split}$ (45)
@@
IV-F ALGORITHM ANALYSIS
@@ IV-F2 Complexity Anlysis
In Algorithm 1, we solve the subproblem SP1 using the binary cut-and-branch method [2013MunozJoint], and solve other subproblems using the interior point method [2010LuoSemidefinite]. Thus, the computational complexity of the four subproblems are $\mathcal{O}\left(MKN\log(MKN)\right)$, $\mathcal{O}\left((MN)^{3.5}\log(1/\epsilon)\right)$, $\mathcal{O}\left((MKN)^{3.5}\log(1/\epsilon)\right)$, $\mathcal{O}\left((MKN)^{3.5}\log(1/\epsilon)\right)$, respectively, where $\epsilon$ represents the tolerance. Consequently, the complexity of Algorithm 1 is $\mathcal{O}(\mathcal{I}(((MN)^{3.5}$ $(MKN)^{3.5}+(MKN)^{3.5})\log(1/\epsilon)+MKN\log(MKN)))$, where $\mathcal{I}$ denotes the iteration number.
@@ V NUMERICAL RESULTS
In the simulations, we consider $K=2$ UAVs fly from specified initial positions to specified final positions to provide computing services for the MUs. We consider $M=8$ MUs that are randomly distributed in a square area of $1000\times 1000\ \rm{m}^{2}$. All UAVs fly at a height of 100 $\rm{m}$. Unless otherwise stated, the remaining simulation parameters are summarized in Table II [Chen2022Energy].
TABLE II: Simulation Parameters.
@@
Number of time slot, NN 40
@@
Channel power gain between the UAVs and the MUs, $\beta_{0}$ -60 dBm
Noise power at the UAV, $\sigma_{\rm{UAV}}$ -170 dBm
@@
Noise power at the LEO satellite, $\sigma_{\rm{LEO}}$ -170 dBm
@@
To evaluate the performance and efficiency of the proposed scheme, we compare it with four baseline schemes, which are given in the following.
@@
    LS Offloading Scheme: Only the LEO satellite provides computing services for the MUs.
@@
Fig. 5 and Fig. 6 show that, as the number of MUs increases, our proposed scheme exhibits a more significant increase in total energy efficiency compared to other four schemes, while the increase in energy consumption is relatively smaller. The reason can be explained as follows. As the number of MUs grows, the demand for the UAV services increases, and the various schemes exhibit different levels of adaptability to these changes. Specifically, the single UAV scheme deploys only a single UAV, which makes it challenging to ensure high-quality communication services for every MU when a large number of MUs require for the services. The fixed trajectory scheme lacks flexibility, especially when there are multiple UAVs. This means that in scenarios with the uneven MU distribution or dynamic demand changes, the scheme struggles to optimize service quality by adjusting the UAV positions. The fixed allocation scheme employs a fixed task allocation decision, which means it cannot adjust the allocation of tasks based on real-time network conditions. In contrast, our scheme dynamically reallocates tasks according to the current network state, ensuring higher service quality and better user experience. The LS offloading scheme relies solely on local computing resources and the LEO satellite assisted computing without considering the role of the UAVs. Although this approach can provide a certain level of service, the absence of the UAV support makes it difficult to achieve ideal computational quality and efficiency in scenarios with a large-scale MU distribution. Therefore, under the same number of MUs, our proposed scheme demonstrates the best performance. Specifically, from the numerical results, when the number of the MUs is $16$, our proposed scheme outperforms the single UAV scheme, fixed trajectory scheme, fixed allocation scheme, and LS offloading scheme by $17.53\%$, $22.55\%$, $48.27\%$, and $133.90\%$, respectively, in terms of the overall computational energy efficiency.
@@
Fig. 7 and Fig. 8 show that, as the amount of task data that the MUs need to process per time slot increases, the total energy efficiency of all schemes decreases, and the required energy consumption also increases. Notably, in both key metrics of total energy efficiency and energy consumption, our proposed algorithm outperforms other four algorithms. The reason is that as the amount of task data increases, the MUs tend to offload more tasks to the UAVs and the LEO satellite for remote computing, leading to an increase in communication energy consumption, which in turn reduces the total energy efficiency of all schemes. Moreover, the LS offloading scheme experiences the most significant increase in total energy consumption, and the rate of increase is notably faster than that of other four schemes. The reason is that, in the LS offloading scheme, the MUs do not offload tasks to the UAVs.
@@
Fig. 9 presents the total computational energy efficiency versus the communication bandwidth between the UAVs and the MUs. It can be observed that the total energy efficiency significantly increases with the growth of the communication bandwidth between the MUs and the UAVs for the proposed scheme, single UAV scheme, and fixed trajectory scheme, whereas the increase is relatively slower for fixed allocation scheme, and remains unchanged for LS offloading scheme. The reason is that the increase in communication bandwidth enhances the task transmission rate between the MUs and the UAVs, allowing the UAVs to assist in processing a larger volume of tasks within each time slot $\tau$. Specifically, for the proposed scheme, single UAV scheme, and fixed trajectory scheme, the increase in communication bandwidth means that the MUs can offload tasks to the UAVs more quickly, improving the efficiency of task processing and reducing the burden of local computation, thus enhancing overall energy efficiency. However, in the fixed allocation scheme, since the proportions of task offloading to the UAVs is fixed, the growth in communication bandwidth can only reduce the energy consumption associated with the UAV computing, but does not enable the MUs to offload more tasks to the UAV, thereby limiting its potential for improvement in energy efficiency. For the LS offloading scheme, since there is no UAV involved, the increase in communication bandwidth has no effect, and the total energy efficiency remains a constant.
Fig. 10 presents the total energy consumption versus the communication bandwidth between the UAVs and the MUs. The proposed scheme, single UAV scheme, and fixed trajectory scheme demonstrate a significant reduction in total energy consumption as the communication bandwidth increases, whereas the reduction is relatively slower for the fixed allocation scheme, and remains unchanged for the LS offloading scheme. The reason is that the increase in communication bandwidth significantly enhances the communication rate between the MUs and the UAVs, thereby increasing the proportion of tasks offloaded from the MUs to the UAVs, which is shown in Fig. 11. Specifically, when the communication bandwidth increases from $1$ MHz to $2$ MHz, the average proportion of tasks offloaded to the UAVs rises from $49\%$ to $65\%$, while the average proportion of tasks processed locally decreases from $31\%$ to $24\%$ in the proposed scheme. Since the MUs are equipped with low energy-efficient processors, offloading more tasks to the UAVs for processing can effectively lower the total energy consumption of the system.